# Deep Reinforcement Learning for Traveling Purchaser Problems

### v5

@@ 7.4 Results on TPPLIB Benchmark Instances
Table 4: Results on TPPLIB benchmark instances.
@@
Then, we evaluate our approach on the TPPLIB benchmark, by directly running the policy network (trained on synthetic instances) on the benchmark instances. The TPPLIB benchmark contains 5 instances for each instance distribution as a category. The results are summarized in Table 4. The first block of the table (prefixed with “EEuclideo”) reports the results on U-TPP instances, and the following three blocks (prefixed with “CapEuclideo”) report the results on R-TPP instances with $\lambda=0.99,0.95,0.9$, respectively. Each row presents the average result on a category of 5 benchmark instances. For example, “CapEuclideo.50.100.95” denotes the R-TPP instances with $|M|=50,|K|=100,\lambda=0.95$. We report 1) the average objective value, 2) the optimality gap, and 3) the runtime (in seconds) for each category of the TPPLIB benchmark.
The results demonstrate that the policy network, trained on synthetic instances, can be effectively applied to TPPLIB benchmark instances from the same instance distribution. Our DRL-based approach outperforms the baseline heuristics by a large margin in terms of the optimality gap, especially for large-sized R-TPP instances. In detail, the average optimality gaps of solutions produced using our RL+TRH method are consistently within 6% on each category of the TPPLIB benchmark in the experiment, yielding a reduction ranging from 40% to 90% compared to the baseline heuristics. The comparison of optimality gaps is illustrated in Figure 6. Besides, similar to the results on synthetic instances, our DRL-based approach also shows a significant advantage over the baseline heuristics in terms of the runtime.
@@ 7.5 Generalization Performance on Large-Sized Instances
Finally, we examine the zero-shot generalization performance of our approach. We use the meta-learning strategy to train the policy network on TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$, and then directly apply the meta policy network to solve larger-sized problem instances that are never seen in the training stage (without any fine-tuning). The results are summarized in Table 5. The first block of Table 5 presents the results on U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$, $(200,200)$, and $(300,300)$, and the second block presents the results on R-TPP instances with $|M|=150,|K|=150$ and $\lambda=0.99,0.95,0.9$.
It is demonstrated that our DRL-based approach still shows an advantage over the baseline heuristics when generalizing to U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$ and $(200,200)$, while the largest instances for training are only of size $\left(|M|,|K|\right)=(100,100)$. However, we observe a drop in the zero-shot generalization performance on U-TPP instances of size over $\left(|M|,|K|\right)=(300,300)$ and R-TPP instances of size over $\left(|M|,|K|\right)=(150,150)$. In future work, we shall further explore the generalization and more efficient transfer learning approaches for larger-sized TPP instances.
@@
Table 5: Results on larger-sized instances.
@@ 8 Conclusion
In this paper, we have presented a novel DRL-based approach for solving TPPs, which exploits the idea of “solve separately, learn globally”. Namely, we break the solution task into two separate stages at the operating level: route construction and purchase planning, while learning a policy network towards optimizing the global solution objective. Built on this framework, we proposed a bipartite graph representation for TPPs and designed a policy network that effectively extracts information from the bipartite graph for route construction. Moreover, we introduced a meta-learning strategy, which significantly enhances the training stability and efficiency on large-sized instances and improves the generalization ability. Experimental results demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics in both solution quality and runtime, while also showing good generalization performance. Future works shall further explore better and more efficient generalization to larger-sized TPP instances and a more generic framework to address other TPP variants, such as multi-vehicle TPP, dynamic TPP, etc.
Acknowledgments
This work was supported by the National Natural Science Foundation of China [Grants 61936009 and 42327901], and was supported by the Research and Development Project of Lenovo Group Ltd. and the Research and Development Project of CRSC Research & Design Institute Group Co., Ltd.

### v6

@@ 7.4 Results on TPPLIB Benchmark Instances
We next evaluate our approach on the TPPLIB benchmark, by directly running the policy network (trained on synthetic instances) on the benchmark instances. The TPPLIB benchmark contains 5 instances for each instance distribution as a category. The results are summarized in Table 4. The first block of the table (labeled with “EEuclideo”) reports the results on U-TPP instances, and the following three blocks (labeled with “CapEuclideo”) report the results on R-TPP instances with $\lambda=0.99,0.95,0.9$, respectively. Each row presents the average result for a category of 5 benchmark instances. For example, “CapEuclideo.50.100.95” refers to R-TPP instances with $|M|=50,|K|=100,\lambda=0.95$. We report 1) the average objective value, 2) the average optimality gap, and 3) the average runtime (in seconds) for each category of the TPPLIB benchmark.
Table 4: Results on TPPLIB benchmark instances
@@
Average 2912 187 3257 14.66% 0.212 3274 14.94% 0.451 3068 5.51% 0.033 2995 2.83% 0.057
The results demonstrate that the policy network, trained on synthetic instances, can be effectively applied to TPPLIB benchmark instances from the same instance distribution. Our DRL-based approach outperforms the baseline heuristics by a large margin in terms of the optimality gap, especially for large-sized R-TPP instances. In detail, the average optimality gaps of solutions produced using our RL+TRH method are consistently within 6% on each category of the TPPLIB benchmark in the experiment, yielding a reduction ranging from 40% to 90% compared to the baseline heuristics. The comparison of optimality gaps is illustrated in Figure 6. Besides, similar to the results on synthetic instances, our DRL-based approach also shows a significant advantage in runtime.
@@ 7.5 Generalization Performance on Large-Sized Instances
Furthermore, we validate the zero-shot generalization performance of our DRL-based approach. We employ the meta-learning strategy to train the policy network on TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$, and then directly apply the meta policy network to solve larger-sized problem instances that are never seen in the training stage, without any fine-tuning. The results are summarized in Table 5. The first block of Table 5 presents the results on U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$, $(200,200)$, and $(300,300)$, and the second block presents the results on R-TPP instances with $|M|=150,|K|=150$ and $\lambda=0.99,0.95,0.9$.
It is demonstrated that our DRL-based approach still shows an advantage over the baseline heuristics when generalizing to U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$ and $(200,200)$, while the largest instances for training are only of size $\left(|M|,|K|\right)=(100,100)$. However, we observe a drop in the zero-shot generalization performance on U-TPP instances of size over $\left(|M|,|K|\right)=(300,300)$ and R-TPP instances of size over $\left(|M|,|K|\right)=(150,150)$. In future work, we shall further explore the generalization and more efficient learning techniques for larger-sized TPP instances.
Table 5: Results on larger-sized instances
@@
In addition, to get a better understanding of the contributions of components in our framework, we conduct a series of ablation studies. The results are presented in B. We also present extended results that include comparisons with state-of-the-art neural VRP solvers and a validation of the cross-problem compatibility with other VRP variants in C to further highlight the effectiveness and compatibility of our approach.
@@ 8 Conclusion
In this paper, we have presented a novel DRL-based approach for solving TPPs, which exploits the idea of “solve separately, learn globally”. Namely, we break the solution task into two separate stages at the operational level: route construction and purchase planning, while learning a policy network towards optimizing the global solution objective. Built on this framework, we proposed a bipartite graph representation for TPPs and designed a policy network that effectively extracts information from the bipartite graph for route construction. Moreover, we introduced a meta-learning strategy, which significantly enhances the training stability and efficiency on large-sized instances and improves the generalization ability. Experimental results demonstrate that our DRL-based approach can significantly outperform well-established TPP heuristics in both solution quality and runtime, while also showing good generalization performance. Future works shall further explore better and more efficient generalization to larger-sized TPP instances and a more generic framework to address other TPP variants, such as multi-vehicle TPP, dynamic TPP, etc.
Appendix A Related Work
Here we review the existing methods for TPPs, DRL-based methods for routing problems, and meta-learning.
A.1. Traveling Purchaser Problem
The TPP is an important combinatorial optimization problem in logistics and manufacturing, with broad applications such as purchasing required raw materials for manufacturing factories (Ramesh, 1981), or scheduling a set of jobs for certain machines (Burstall, 1966), etc. Notably, the TSP is a special case of TPPs (Manerba et al., 2017).
Existing methods for solving TPPs can be categorized into exact methods and heuristics (Manerba et al., 2017). The first exact method, introduced in (Buzacott and Dutta, 1971), solved U-TPP using dynamic programming. Subsequently, Singh and van Oudheusden (1997) developed a branch-and-bound algorithm, capable of solving TPP instances with up to 20 markets and 100 products, or 25 markets and 50 products. Laporte et al. (2003) further enhanced this approach through a branch-and-cut algorithm that exploits dynamic generation of variables and separation of constraints. Riera-Ledesma and González (2006) extended this methodology to general TPPs. Despite a significant advance in the size of TPPs solved to optimality, their computational cost remains unaffordable in real-time applications.
Alternatively, various heuristic methods have emerged to generate high-quality solutions within a reasonable time. GSH (Golden et al., 1981), the first constructive heuristic for TPPs, starts from the depot and iteratively inserts the unvisited market offering the largest cost saving. Ong (1982) introduced TRH, a reduction-based heuristic that starts with an initial route containing a subset of markets, and iteratively drops the market yielding the largest cost reduction. The CAH (Pearn and Chien, 1998) adopts a product-focused approach, constructing a least-cost solution for the first product and, in subsequent iterations, adding each product to the solution in a least-cost manner. In practice, GSH and CAH are usually followed by TRH and a TSP solver to remove redundant markets and optimize market sequencing for further improvements. However, these heuristics rely heavily on substantial specialized expertise and knowledge, and their performance remains limited. Additional works, such as local search methods, explore improved solutions at the expense of extra computational time (Teeninga and Volgenant, 2004; Riera-Ledesma and González, 2005), but, overall, there was no significant breakthrough in heuristics for TPPs.
A.2. DRL for Routing Problems
DRL-based methods for routing problems can be classified into two categories: constructive methods and improvement methods. Constructive methods focus on learning policies to construct solutions step-by-step. Vinyals et al. (2015) pioneered a sequence-to-sequence model, the Pointer Network (Ptr-Net), to solve the TSP in an autoregressive manner. Later, with the development of self-attention mechanism, Kool et al. (2019) introduced an attention-based model (AM), which achieved superior performance in various routing problems and became a milestone in this field. Since then, several AM-variants have been proposed, with most advanced neural solvers built on top of them (Gao et al., 2024; Kwon et al., 2020; Xin et al., 2021). However, due to their limitations in the decision framework, problem representation, and training scheme, these approaches can only handle routing problems with simple structures and cannot be readily extended to more complex problems such as TPPs. Notably, while recent research has explored general models for different VRP variants (Lin et al., 2024; Zhou et al., 2024), these efforts still remain confined to combinations of straightforward VRP configurations.
As another line of research, improvement methods leverage DRL to iteratively refine an initial solution through local search (Chen and Tian, 2019; Wang et al., 2025). It is worth mentioning that our approach is compatible with such improvement methods, and could potentially be integrated for further improvement in applications with larger computational budgets.
A.3. Meta-Learning
Meta-learning, often referred to as ”learning to learn”, is a machine learning paradigm designed to enhance the adaptability and efficiency of learning algorithms across new tasks or data distributions by leveraging prior training experience (Barman et al., 2025). This idea of meta-learning traces back decades ago, aiming to enable systems to adapt to new tasks with minimal data by leveraging prior knowledge, similar to how humans learn new skills based on past experiences (Schmidhuber, 1987). A significant milestone in meta-learning was the introduction of model-agnostic meta-learning (MAML) (Finn et al., 2017b), which learns the initial parameters of neural networks such that it can be efficiently fine-tuned with a few gradient steps for fast adaptation to new tasks. Follow-up works have focused on improving the efficiency and robustness (Goldblum et al., 2020), and exploring other application domains such as RL (Beck et al., 2023). To date, the applications of meta-learning have expanded to complex tasks such as robotics (Schoettler et al., 2020), optimization (Qiu et al., 2022), and multi-agent systems (Munir et al., 2021). For a comprehensive overview, readers are referred to the survey by (Barman et al., 2025).
Appendix B Ablation Study
To get a better understanding of the contributions of components in our framework, we conduct a series of ablation studies. Following the organization of this paper, we study the contributions of each key component, including the bipartite graph representation, the policy network architecture, and the meta-learning strategy. The experimental results are presented in Table 6, where we compare the performance of RL-E2E to exclude the effect of the post-optimization procedure. Besides, as the differences in runtime are very marginal, we focus on the objective value for comparative analysis.
B.1. Bipartite Graph Representation
We evaluate the contributions of our bipartite graph representation by comparing it with the 1) complete graph and 2) k-NN graph representations (entries “Cplt.” and “k-NN” in Table 6). The node features are defined as described in Section 4, and the k-NN graphs are adapted using attention masks in the encoder, with neighbor selection following (Joshi et al., 2019). The results show that, while the complete and k-NN graphs encode product supply information through node features, our bipartite graph representation facilitates the capture of high-level information for both markets and products through edges connecting the market nodes and product nodes, thus leading to better solutions. Notably, the k-NN graph representation exhibited a significant performance drop, contrasting with its effectiveness in TSP. This is because the route construction for TPPs should take into account not only the spatial relations between markets but also their purchasing interdependencies. Consequently, k-NN graphs, which emphasize spatial proximity, can potentially misguide the route construction and result in poor performance.
Table 6: Ablation study on representation and network architecture
Instance  Full  Model Repre. Emb. Module Dec. Context
$|M|$ $|K|$ $\lambda$ Cplt.1 k-NN2 IEM3 ME4 LSTM5 GE6 DC7 RC8
50 50 / 1897 1941 1922 2985 1986 1911 1928 1933 1976
50 100 / 2542 2570 3521 3923 2712 2570 2583 2597 2630
100 50 / 1563 1599 1574 2897 1689 1579 1612 1667 1743
100 100 / 2111 2159 2834 4010 2296 2125 2165 2205 2338
50 50 0.99 2032 2084 2528 3085 2102 2052 2048 2048 2197
50 100 0.99 2567 2865 3616 3865 2762 2581 2595 2611 2883
100 50 0.99 1753 1799 2206 3067 1962 1778 1818 1888 2161
100 100 0.99 2302 2434 3348 4111 2607 2338 2351 2450 2792
50 50 0.95 2645 2680 3501 3521 2857 2663 2714 2668 3065
50 100 0.95 3457 3531 4930 4411 3645 3497 3488 3485 4171
100 50 0.95 3027 3106 3536 4118 3451 3098 3138 3087 3711
100 100 0.95 3993 4165 5629 5240 4461 4058 4073 4079 4884
50 50 0.9 3704 3763 4900 4552 3975 3744 3777 3769 4235
50 100 0.9 4927 5047 6563 5737 5184 4985 4951 4967 5658
100 50 0.9 4956 5110 5266 6016 5580 5047 5121 5036 6034
100 100 0.9 6672 6869 7715 7823 7531 6869 6899 6838 8209
  • 1
    complete graph 2 k-NN graph 3 w/o input embedding module 4 w/o market encoder
  • 5
    w/o LSTM module 6 w/o global embedding 7 w/o demand context 8 w/o route context
B.2. Policy Network Architecture
We study the contributions of the key components in our policy network, focusing on the embedding modules and the decoding context. Specifically, for the embedding modules, we evaluate the removal of 1) the input embedding module (IEM), 2) the market encoder (ME), and 3) the LSTM for current route embedding in the decoder. For the decoding context, we evaluate the removal of 1) the global embedding (GE), 2) the demand context (DC), and 3) the route context (RC). The results are presented in Table 6 under categories “Emb. Module” and “Dec. Context”.
The results indicate that all three embedding modules contribute to overall performance, with the input embedding module being the most critical. Removing the input embedding module leads to a significant performance drop, as its removal separates the market locations from the product supply information in the input. This underscores the necessity of jointly considering spatial and supply information for TPPs. The market encoder becomes increasingly important with larger problem sizes, particularly for instances with more markets, owing to its capacity to extract deeper relational information between markets. Furthermore, while LSTM-based embeddings for the current partial route are not commonly employed in AM-based neural solvers (Kool et al., 2019), we find that this component enhances solution quality for TPPs, since all visited markets can influence subsequent decisions in solving the TPP.
Interestingly, the removal of global embedding has only a marginal impact on the solution quality. This suggests that the first one-to-many attention layer in the decoder can also serve as a mechanism for aggregating global information. Moreover, as discussed in Section 3.2.1, explicitly providing the demand context to the decoder also contributes to improved performance, especially when accurately meeting the demands plays a more important role in route construction. Finally, it is noteworthy that removing the route context, where our model becomes nearly non-autoregressive, still yields reasonable solutions. This suggests our policy network holds strong potential for non-autoregressive modeling—a parallelizable framework considered advantageous for scaling to extremely large instances (Fu et al., 2021; Sun and Yang, 2023). This finding opens up possibilities for further improvements in scalability in future work.
B.3. Meta-Learning Strategy
The necessity and effectiveness of our meta-learning strategy for training stability have been discussed in Section 7.2. Here we focus on its impact on the zero-shot generalization ability, by comparing it against two commonly-used learning strategies: 1) transfer learning (TL) (Montanez-Barrera et al., 2024), where a policy network pre-trained on small instances is used as initialization for training on large instances, and 2) curriculum learning (CL) (Zhou et al., 2023), which progressively increases the size of training instances, starting from smaller ones and gradually scaling to larger ones. The comparisons of zero-shot generalization performance are reported in Table 7, demonstrating the advantage of our meta-learning strategy in generalizing to previously unseen, larger-sized instances.
Table 7: Comparisons of zero-shot generalization
Instance Heuristic Meta TL CL
$|M|$ $|K|$ $\lambda$   GSH +TRH   CAH +TRH E2E1 TRH2 E2E TRH E2E TRH
150 150 / 2672 2460 2514 2380 2745 2458 2648 2398
200 200 / 2885 2454 2485 2262 2850 2497 2806 2339
300 300 / 3445 3074 3206 3088 4367 3243 4043 3164
150 150 0.99 2918 3056 2576 2519 2797 2604 2898 2694
150 150 0.95 5642 5884 6133 6014 6316 6174 6454 6224
150 150 0.9 10008 10199 10707 10155 12061 10908 16027 12478
  • 1
    RL-E2E based on meta-learning strategy
  • 2
    RL+TRH based on meta-learning strategy
Appendix C Extended Results
At last, to further highlight the effectiveness and compatibility of our approach, we present extended results that include comparisons with state-of-the-art neural VRP solvers, as well as a validation of the cross-problem compatibility of our approach and network architecture with other VRP variants.
C.1. Comparisons with Other Neural Solvers
We compare our approach with several neural solvers, including the widely-recognized Ptr-Net (Vinyals et al., 2015) and AM (Kool et al., 2019), as well as the state-of-the-art unified neural VRP solver, MVMoE (Zhou et al., 2024). Since the baseline neural VRP solvers are not readily applicable to TPPs, we adapt their node features as described in Section 4, and employ transfer learning to mitigate potential training collapse on large-sized instances. The results are presented in Table 8. Our approach consistently outperforms the above neural solvers across all instance sets.
Beyond superior solution quality, our approach offers a significant advantage in scalability. The baseline neural solvers, limited by their problem representations, require separate models and training procedures for TPP instances with varying numbers of products. Even MVMoE, despite introducing a unified policy network with feature padding to multiple VRP variants, remains restricted by a fixed-dimensional feature representation, limiting its flexibility across TPP instances of varying sizes. This limitation not only hinders them from learning generalized knowledge but also significantly increases the training and storage overhead in practice. In contrast, our bipartite graph representation enables a size-agnostic model, allowing us to train a single, generalizable policy network through meta-learning that effectively adapts across varied instance sizes and distributions, significantly enhancing its flexibility and efficiency for real-world applications.
Table 8: Comparisons with other neural solvers
Instance Ptr-Net AM MVMoE Ours
$|M|$ $|K|$ $\lambda$ E2E TRH E2E TRH E2E TRH E2E TRH
50 50 / 2427 2056 2009 1906 2210 1976 1897 1857
50 100 / 3357 2677 2856 2525 2813 2507 2542 2446
100 50 / 2284 2027 1643 1556 1938 1727 1563 1524
100 100 / 3114 2394 2404 2148 2476 2186 2111 2044
50 50 0.99 2997 2430 2133 2005 2272 2049 2032 1954
50 100 0.99 3891 2821 3093 2650 2817 2546 2567 2466
100 50 0.99 2284 2027 1863 1777 2123 1927 1753 1711
100 100 0.99 3598 2867 2720 2465 2665 2421 2302 2235
50 50 0.95 3281 2925 2843 2678 2922 2723 2645 2594
50 100 0.95 4255 3788 3857 3529 3714 3469 3457 3368
100 50 0.95 3744 3451 3280 3137 3490 3284 3027 2980
100 100 0.95 4996 4523 4400 4162 4465 4218 3993 3920
50 50 0.9 4391 4014 3938 3758 4024 3819 3704 3644
50 100 0.9 5676 5257 5305 5066 5289 5025 4927 4855
100 50 0.9 5864 5488 5282 5134 5623 5313 4956 4900
100 100 0.9 7664 7246 7223 6946 7426 7018 6672 6660
C.2. Compatibility with Other VRP Variants
It is worth mentioning that many VRP variants can be formulated as special cases of the TPP. For example, the TSP is equivalent to the UTPP where each market provides a unique product with zero price. Similarly, the prizes and penalties in routing problems such as the orienteering problem (OP) (Golden et al., 1987) and the prize collecting TSP (PCTSP) (Balas, 1989) can also be represented as products with associated purchasing costs in TPPs. Therefore, our approach is inherently compatible with these VRP variants. To validate this, we evaluate our approach on the TSP, OP, and PCTSP, following the experimental settings in (Kool et al., 2019). The results are summarized in Table 9. It is demonstrated that our approach outperforms AM on most problems, mainly due to our improvements in the embedding module, decoding context, and training scheme. Moreover, our approach could potentially be further enhanced by incorporating training and inference techniques tailored to these specific VRP variants.
Table 9: Compatibility with other VRP variants
Method TSP ($\downarrow$)* OP ($\uparrow$)* PCTSP ($\downarrow$)*
$n$=20 $n$=50 $n$=100 $n$=20 $n$=50 $n$=100 $n$=20 $n$=50 $n$=100
AM 3.85 5.80 8.12 5.19 15.64 31.62 3.18 4.60 6.25
Ours 3.84 5.73 7.97 6.54 16.20 31.49 3.04 4.48 6.11
    $\downarrow$ means a minimization problem, and $\uparrow$ means a maximization problem